I need help on how to solve quadratic equations using square roots. I have two problems I need help with:
1.) (2x+1)^2=25

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

1.) (2x+1)^2=25
------
2x+1 = 5
or
2x+1 = -5
-------------
2.) (x-7)^2=1
x-7 = 1
or
x-7 = -1
===============

Answer by richwmiller(17219)

(2x+1)^2=25
2x+1=5
2x=4
x=2
2x+1=-5
2x=-6
x=-3